# Reducing Biases in Record Matching Through Scores Calibration

### v2

@@ 17       $\text{pos}_{g}\leftarrow\text{pos}_{g}+1$;
Following Example 2, consider a dataset $D$ containing 15 pairs with scores as shown in Figure 3. Using meanshift, the decision threshold is set to $\gamma=0.57$, and the query pair $p$ is predicted as 0 (nonmatch). As a result, the calibration process is applied only to pairs predicted as nonmatches (label 0). As shown in Figure 3, the matched pairs are filtered out, and the remaining pairs are calibrated as in Algorithm 1. This leads to sorted scores in descending order $\text{scores}_{b}=[0.37,0.35,0.31,0.25,0.22,0.18]$ and $\text{scores}_{a}=[0.46,0.39,0.28]$. Next, we calculate $\text{size}=9$ (the total number of non-matching pairs in $D$ to be used instead of $|D|$ for computing $\alpha$), with $|\text{scores}_{a}|=3$ and $|\text{scores}_{b}|=6$. The positions are $\text{pos}_{a}=2$ and $\text{pos}_{b}=3$ with quantile $q=0.5$. We also compute $\alpha=|\text{scores}_{a}|/\text{size}=3/9=0.33$. The calibrated score is calculated as $\hat{s}(p)$ = $0.33\times 0.39$ + $0.67\times 0.31\approx 0.33$ Therefore, the CondCalibrate algorithm returns a calibrated score of $0.33$ for the query pair $p$. $\blacksquare$
@@
VI Experiments
In our experiments, we first assess bias in existing matching methods using our score bias metric. Second, we show that this metric captures bias in matching tasks more effectively than traditional binary measures (Section VI-B1). Third, we evaluate our calibration algorithms for their ability to reduce score bias while maintaining low risk. Bias reduction results are presented in Sections VI-B2 and VI-B3, with risk analysis in Section VI-B4. Finally, we examine the impact of different input dataset $D$ selections in Section VI-B5.
VI-A Experimental Setup
@@
VI-A1 Datasets
We used record matching benchmarks [39]: Beer (BEER), Walmart-Amazon (WAL-AMZ), Amazon-Google (AMZ-GOO), Fodors-Zagat (FOD-ZAG), iTunes-Amazon (ITU-AMZ), DBLP-GoogleScholar (DBLP-GOO), and DBLP-ACM (DBLP-ACM). Table II summarizes each dataset statistics. None of these datasets explicitly denotes a sensitive attribute, we follow prior work [7, 8, 40] to define minority and majority based on the following criteria:
@@ VI-A1 Datasets
Dataset #Attr. Dataset Size EqSens. Attr.
@@
TABLE II: Datasets and their characteristics. Dataset Size shows training and test set sizes. Eqcolumns represent the percentage of equivalence pairs in the training and test data.
VI-A2 Record Matching Methods
@@
We selected five high-performing deep learning record-matching methods with diverse architectures to demonstrate our calibration method’s versatility. DeepMatch combines tokenized inputs, pre-trained embeddings, and metadata, achieving strong performance on structured data [39]. HierGAT employs a Hierarchical Graph Attention Transformer for contextual embeddings [41]. DITTO enhances pre-trained language models with text summarization, data augmentation, and domain expertise [42]. EMTransform uses transformers to model relationships across attributes [43]. HierMatch builds on DeepMatch with cross-attribute token alignment and attribute-aware attention [9]. EMTransform is based on pre-trained transformers and language models  [43]. Each model is trained for 100 epochs, with early stopping if the F1 score fails to improve for 15 epochs. The data is split 75% for training and 25% for validation. After training, the model generates scores for the test set, which feed into our calibration algorithms.
VI-B Experimental Results
@@
We now present our results. For simplicity, we denote matching score biases $s$ as DP, EO, and EOD, corresponding to $\textit{bias}(s,\text{PR})$, $\textit{bias}(s,\text{TPR})$, and $\textit{bias}(s,\text{TPR})+\textit{bias}(s,\text{FPR})$, respectively. We also use $\text{DP}_{\theta}$ and $\text{EOD}_{\theta}$ to represent the demographic disparity and equalized odds difference for a matcher with score function $s$ and threshold $\theta$ (e.g., $\text{DP}_{0.5}$ for $\theta=0.5$ indicates the conventional demographic disparity).
VI-B1 Biases in Matching Scores
@@
Here, we analyze the matching methods and their associated biases. Table III summarizes these biases across datasets and matching models. Each metric is reported for three different thresholds, alongside the bias metric for score values as defined in this work. Due to space limitations, we present a subset of model-dataset combinations here; the complete results are available in our repository [38]. An initial look at the table shows that traditional fairness metrics can vary significantly across thresholds. For example, the biases highlighted in blue indicate a notable disparity with $\text{EO}_{0.1}=14$, but much lower biases for $\text{EO}_{0.5}=1.35$ and $\text{EO}_{0.95}=3.38$. Similarly, the biases highlighted in green show minimal bias for $\text{DP}_{0.1}=0.05$ but much higher biases for $\text{DP}_{0.5}=4.85$ and $\text{DP}_{0.95}=3.49$. Similar trends are evident for EOD across methods and datasets. The score bias metrics shown with DP, EO, and EOD provide an aggregate view of biases across all thresholds, capturing overall bias effectively. The results in the table also reveal differences in biases across methods and datasets. In most cases, substantial biases are present, indicating a clear need for debiasing methods.
@@
TABLE III: We evaluated the distributional disparity across models and datasets. Our metrics were compared against traditional fairness measures at thresholds of 0.1, 0.5 and 0.95.
To further demonstrate the importance of biases for scores (Distributional DP or EO), we analyze two examples from Table III in more detail. These examples are highlighted in Table III and shown in Figure 4. Figure 4a illustrates $\text{DP}_{\theta}$ across varying thresholds for HierMatch on DBLP-ACM, while Figure 4b shows $\text{EO}_{\theta}$ for DeepMatch on AMZ-GOO. As seen in these figures, a model may appear fair at certain thresholds (with narrow color shade widths) based on traditional fairness metrics. However, at thresholds very close to these fair points, the model often displays significant bias toward one group, complicating fairness evaluation. In both figures, the entire colored area effectively represents the distributional disparity.
@@
VI-B2 Bias Reduction through Calibration
Here, we analyze the performance of Calib. For input dataset $D$ , we used the entire test dataset and update query points within this dataset. Figure 5 presents selected results, showing a notable reduction in score bias DP before and after calibration across models and datasets. This shows the effectiveness of the algorithm in minimizing DP and, in many cases, nearly eliminating bias. Due to space limitations, we display a subset of results; similar findings for other combinations are available in our repository. Overall, this calibration approach effectively reduces DP across all tested algorithms and datasets.
@@ VI-B2 Bias Reduction through Calibration
VI-B3 Bias Reduction in Conditional Calibration
Figure 6 also shows the results for C-Calib. As evident in this Figure, there is a notable improvement in reducing both EO and EOD across all models and datasets compared to Calib. The conditional approach is more effective in reducing these fairness metrics, which aligns with its design of adjusting score distributions conditioned on estimated labels. For instance, C-Calib in AMZ-GOO dataset (Figure 6a), reduced EOD more effectively across all models compared to Calib. Specifically, HierGAT’s EOD for AMZ-GOO was increases to 30.92% from its initial value of 19.50% by Calib, however C-Calib reduced it to 7.39%. Similarly, in BEER dataset (Figure 6b), the EO for DITTO on HierGAT are significantly reduced to near-zero values post C-Calib, indicating a strong improvement from Calib. Notably, for DITTO on BEER dataset in Figure 6b, the initial EO was 0.30%, which is already minimal, but the C-Calib further reduced it to 0.13%, highlighting the effectiveness of this method. Across other datasets (available in our repository), this pattern of reduced disparity by C-Calib continues, showing that the conditional approach better aligns score distributions across demographic groups by estimating labels.
@@ VI-B3 Bias Reduction in Conditional Calibration
(a) $\Delta$EO using Calib
(b) $\Delta$EOD using Calib
(c) $\Delta$EO using C-Calib
(d) $\Delta$EOD using C-Calib
@@
VI-B4 Risk of Calibration
To conclude our analysis, we examine the impact of Calib and C-Calib on the risk. Risk, as defined in Definition 2, refers to the deviation of calibrated scores from the initial scores. We evaluate this by comparing the accuracy of the matching scores before and after calibration using the AUC metric. Any decrease in AUC indicates a departure from the initial scores that provided a higher AUC. The focus on AUC, rather than other metrics such as F1-Score or precision, is intentional, as these metrics are threshold-dependent. This choice aligns with the goal of this work to provide a threshold-independent analysis.
Table IV shows the effect of Calib and C-Calib on AUC for AMZ-GOO and DBLP-GOO datasets across matching models. For clarity, only the AUC decreases after calibration are reported. In all cases, both algorithms have minimal AUC reductions, and similar results are observed universally, however only a subset of results is reported due to space constraints. In AMZ-GOO dataset, AUC decrease post-calibration using Calib at most 1.01%. Note that a maximum reduction of 1.01% in AUC results in almost completely removing DP bias from an initially significant value. Similarly, in DBLP-GOO dataset, there is almost no change in AUC after calibration using Calib across different matchers, indicating that Calib effectively reduces DP without compromising accuracy.

### v3

@@ 5.1 C-Calib Algorithm
Following Example 2, consider a dataset $D$ containing 15 pairs with scores as shown in Figure 3. Using meanshift, the decision threshold is set to $\gamma=0.57$, and the query pair $p$ is predicted as 0 (nonmatch). As a result, the calibration process is applied only to pairs predicted as nonmatches (label 0). As shown in Figure 3, the matched pairs are filtered out, and the remaining pairs are calibrated as in Algorithm 1. This leads to sorted scores in descending order $\text{scores}_{b}=[0.37,0.35,0.31,0.25,0.22,0.18]$ and $\text{scores}_{a}=[0.46,0.39,0.28]$. Next, we calculate $\text{size}=9$ (the total number of non-matching pairs in $D$ to be used instead of $|D|$ for computing $\alpha$), with $|\text{scores}_{a}|=3$ and $|\text{scores}_{b}|=6$. The positions are $\text{pos}_{a}=2$ and $\text{pos}_{b}=3$ with quantile $q=0.5$. We also compute $\alpha=|\text{scores}_{a}|/\text{size}=3/9=0.33$. The calibrated score is calculated as $\hat{s}(p)$ = $0.33\times 0.39$ + $0.67\times 0.31\approx 0.33.$ Therefore, the CondCalibrate algorithm returns a calibrated score of $0.33$ for the query pair $p$. $\blacksquare$
@@
5.2 Theoretical Analysis of C-Calib
Before presenting the theoretical guarantees for C-Calib, we introduce several quantities that describe the interaction between score calibration and label stratification. Let
$\varepsilon^{\star}\;=\;\inf_{\theta\in[0,1]}\Pr\!\bigl(Y\neq\mathbf{1}\{s(X)\geq\theta\}\bigr)$
be the irreducible thresholding error of the original score $s$. When the class-conditional score distributions overlap, no threshold perfectly separates Y=1Y{=}1 from Y=0Y{=}0, and thus $\varepsilon^{\star}>0$; when they are separable, $\varepsilon^{\star}=0$.
Relatedly, several works study how fairness guarantees degrade when protected information is only available through noisy measurements or proxy signals (e.g., noisy protected attributes) [5, 26]. These results are complementary to our setting: they typically do not analyze OT/Wasserstein-barycenter score calibration, whereas our focus is on how conditioning distributional alignment on an estimated label $\hat{Y}$ affects EO/EOD.
C-Calib uses an unsupervised threshold $\gamma$ (e.g., estimated by mean-shift) to form predicted labels $\hat{Y}=\mathbf{1}\{s(X)\geq\gamma\}$. The realized mislabeling rate
$\varepsilon_{n}\;=\;\Pr(\hat{Y}\neq Y)$
depends on the sample size $n$ through the estimation of $\gamma$. It is convenient to write
$\varepsilon_{n}\;=\;\varepsilon^{\star}\,+\,(\varepsilon_{n}-\varepsilon^{\star}),$
where the excess term $(\varepsilon_{n}-\varepsilon^{\star})$ captures the estimation error introduced by learning the threshold. Under standard regularity (unique split point, smooth and nonvanishing densities), this term typically shrinks at the parametric rate $O_{p}(n^{-1/2})$.
We refer to $\{Y=1\}$ and $\{Y=0\}$ as the true-label strata (positive and negative), and to $\{\hat{Y}=1\}$ and $\{\hat{Y}=0\}$ as the predicted-label strata. C-Calib applies Calib separately within each predicted stratum: the predicted positive stratum drives the TPR component and the predicted negative stratum drives the FPR component. Consequently, the resulting Equalized Odds behavior depends on two sources of error: (i) finite-sample estimation within each stratum, contributing an $O_{p}(n^{-1/2})$ term, and (ii) the quality of the stratification, governed by $\varepsilon^{\star}$ and $(\varepsilon_{n}-\varepsilon^{\star})$.
Theorem 2.
Let $\hat{s}$ be the score function produced by C-Calib, which (i) estimates a threshold $\gamma$ and labels $\hat{Y}=\mathbf{1}\{s(X)\geq\gamma\}$, and (ii) applies Calib within the predicted strata $\{\hat{Y}=1\}$ and $\{\hat{Y}=0\}$. Let
$n$ $=\min_{y\in\{0,1\},\,g\in\{a,b\}}n_{y,g}$
$\varepsilon_{n}$ $=\Pr(\hat{Y}\neq Y)$
$\varepsilon^{\star}$ $=\inf_{\theta\in[0,1]}\Pr\!\bigl(Y\neq\mathbf{1}\{s(X)\geq\theta\}\bigr).$
Under the same regularity assumptions as for Calib,
$\textit{bias}(\hat{s},\text{TPR})+\textit{bias}(\hat{s},\text{FPR})=O_{p}(n^{-1/2})+O\!\bigl(\varepsilon^{\star}+|\varepsilon_{n}-\varepsilon^{\star}|\bigr)$
$\textit{risk}(s^{\star},\hat{s})=O_{p}(n^{-1/2})+O\!\bigl(\varepsilon^{\star}+|\varepsilon_{n}-\varepsilon^{\star}|\bigr),$
where $s^{\star}$ is an oracle EOD-fair map. In particular, when $\varepsilon^{\star}=0$ (threshold-separable classes) and $\gamma$ is consistent (so $|\varepsilon_{n}-\varepsilon^{\star}|=O_{p}(n^{-1/2})$), both the EOD bias and the risk converge at the parametric rate $O_{p}(n^{-1/2})$.
Proof sketch.
If we could stratify by the true label $Y$, applying Calib within $\{Y=1\}$ and $\{Y=0\}$ would align the group-wise CDFs in each stratum. As in Theorem 1, empirical-process fluctuations of CDFs and quantiles contribute an $O_{p}(n^{-1/2})$ error to both TPR and FPR components, and similarly to the risk.
C-Calib instead stratifies using $\hat{Y}$, so each predicted stratum is a mixture of the two true strata. The mixing proportions are governed by the mislabeling rate $\varepsilon_{n}$, yielding an $O(\varepsilon_{n})$ perturbation of the stratum-wise CDFs. By Lipschitz continuity of quantiles in one dimension, these perturbations propagate linearly through the barycenter map computed by Calib, producing an additional $O(\varepsilon_{n})$ contribution to both EOD components (the TPR and FPR terms) and to the risk.
Finally, writing $\varepsilon_{n}=\varepsilon^{\star}+(\varepsilon_{n}-\varepsilon^{\star})$ separates irreducible overlap and threshold-estimation error, completing the bound. ∎
6 Experiments
In our experiments, we first assess bias in existing matching methods using our score bias metric. Second, we show that this metric captures bias in matching tasks more effectively than traditional binary measures (Section 6.2.1). Third, we evaluate our calibration algorithms for their ability to reduce score bias while maintaining low risk. Bias reduction results are presented in Sections 6.2.2 and 6.2.3, with risk analysis in Section 6.2.4. Finally, we examine the impact of different input dataset $D$ selections in Section 6.2.5.
6.1 Experimental Setup
@@
6.1.1 Datasets
We used record matching benchmarks [29]: Beer (BEER), Walmart-Amazon (WAL-AMZ), Amazon-Google (AMZ-GOO), Fodors-Zagat (FOD-ZAG), iTunes-Amazon (ITU-AMZ), DBLP-GoogleScholar (DBLP-GOO), and DBLP-ACM (DBLP-ACM). Table 2 summarizes statistics for each dataset. None of these datasets explicitly denotes a sensitive attribute.
For DBLP-ACM, we use a gender proxy (female first names appearing in the authors field) to form groups, which serves as an example of a sensitive attribute commonly studied in fairness. For the remaining benchmarks, the group definitions (e.g., genre, manufacturer, keyword-based slices) should be viewed as benchmark-derived subpopulations rather than claims about inherently sensitive attributes. Our goal in using these established slices, following prior work [30, 34, 28], is to (i) evaluate whether calibration can repair matching scores whenever systematic score disparities are observed between groups, and (ii) enable direct comparison with existing record-matching methods and fairness-aware entity matching frameworks that use the same benchmarks and group constructions.
We follow prior work [30, 34, 28] to define minority and majority based on the following criteria:
@@ 6.1.1 Datasets
Dataset #Attr. Dataset Size Eq% Sens. Attr.
@@
Table 2: Datasets and their characteristics. Dataset Size shows training and test set sizes. Eq% columns represent the percentage of equivalence pairs in the training and test data.
6.1.2 Record Matching Methods
@@
We selected five high-performing deep learning record-matching methods with diverse architectures to demonstrate our calibration method’s versatility. DeepMatch combines tokenized inputs, pre-trained embeddings, and metadata, achieving strong performance on structured data [29]. HierGAT employs a Hierarchical Graph Attention Transformer for contextual embeddings [44]. DITTO enhances pre-trained language models with text summarization, data augmentation, and domain expertise [45]. EMTransform uses transformers to model relationships across attributes [4]. HierMatch builds on DeepMatch with cross-attribute token alignment and attribute-aware attention [17]. Each model is trained for 100 epochs, with early stopping if the F1 score fails to improve for 15 epochs. The data are split 75% for training and 25% for validation. After training, the model generates scores for the test set, which feed into our calibration algorithms.
6.2 Experimental Results
@@
We now present our results. For simplicity, we denote matching score biases $s$ as DP, EO, and EOD, corresponding to $\textit{bias}(s,\text{PR})$, $\textit{bias}(s,\text{TPR})$, and $\textit{bias}(s,\text{TPR})+\textit{bias}(s,\text{FPR})$, respectively. We use $\text{DP}_{\theta}$, $\text{EO}_{\theta}$, and $\text{EOD}_{\theta}$ to represent the corresponding threshold-specific (binary) fairness metrics for a matcher with score function $s$ and threshold $\theta$ (e.g., $\text{DP}_{0.5}$ for $\theta=0.5$).
6.2.1 Biases in Matching Scores
@@
Here, we analyze the matching methods and their associated biases. Table 3 summarizes these biases across datasets and matching models. All bias values are reported as percentages (%). Each metric is reported for three different thresholds, alongside the bias metric for score values as defined in this work. Due to space limitations, we present a subset of model-dataset combinations here; the complete results are available in our repository [3]. An initial look at the table shows that traditional fairness metrics can vary significantly across thresholds. For example, the biases highlighted in blue indicate a notable disparity with $\text{EO}_{0.1}=14\%$, but much lower biases for $\text{EO}_{0.5}=1.35\%$ and $\text{EO}_{0.95}=3.38\%$. Similarly, the biases highlighted in green show minimal bias for $\text{DP}_{0.1}=0.05\%$ but much higher biases for $\text{DP}_{0.5}=4.85\%$ and $\text{DP}_{0.95}=3.49\%$. Similar trends are evident for EOD across methods and datasets. The score bias metrics shown with DP, EO, and EOD provide an aggregate view of biases across all thresholds, capturing overall bias effectively. The results in the table also reveal differences in biases across methods and datasets. In most cases, substantial biases are present, indicating a clear need for debiasing methods.
@@
Table 3: We evaluated the distributional disparity across models and datasets. Our metrics were compared against traditional fairness measures at thresholds of 0.1, 0.5, and 0.95. All values are reported in %.
To further demonstrate the importance of biases for scores (Distributional DP or EO), we analyze two examples from Table 3 in more detail. These examples are highlighted in Table 3 and shown in Figure 4. Figure 4a illustrates $\text{DP}_{\theta}$ across varying thresholds for HierMatch on DBLP-ACM, while Figure 4b shows $\text{EO}_{\theta}$ for DeepMatch on AMZ-GOO. As seen in these figures, a model may appear fair at certain thresholds (with narrow color shade widths) based on traditional fairness metrics. However, at thresholds very close to these fair points, the model often displays significant bias toward one group, complicating fairness evaluation. In both figures, the entire colored area effectively represents the distributional disparity.
@@
6.2.2 Bias Reduction through Calibration
Here, we analyze the performance of Calib, which requires a set of unlabeled record pairs to estimate score distributions per group. For the input dataset $D$, we use the test inputs themselves (i.e., the set of query pairs) to estimate group-wise score distributions and calibrate their scores. Importantly, this step does not use ground-truth match labels; it only uses the score function outputs on the given inputs. This choice aligns with the intended use case of post-processing a black-box matcher when only the deployment inputs are available.
This experimental protocol follows the standard i.i.d. assumption in machine learning, where training and test inputs are drawn from the same (or sufficiently similar) distribution. When there is substantial distribution shift between the data used to estimate score distributions (i.e., $D$) and the query inputs being calibrated, calibration quality can degrade; studying and addressing such distribution shift is beyond the scope of this work. We partially illustrate this sensitivity by comparing different choices of $D$ in Section 6.2.5.
Figure 5 presents selected results, showing a notable reduction in score bias DP before and after calibration across models and datasets. This shows the effectiveness of the algorithm in minimizing DP and, in many cases, nearly eliminating bias. Due to space limitations, we display a subset of results; similar findings for other combinations are available in our repository. Overall, this calibration approach effectively reduces DP across all tested algorithms and datasets.
6.2.3 Bias Reduction in Conditional Calibration
Figure 6 also shows the results for C-Calib. As evident in this Figure, there is a notable improvement in reducing both EO and EOD across all models and datasets compared to Calib. The conditional approach is more effective in reducing these fairness metrics, which aligns with its design of adjusting score distributions conditioned on estimated labels. For instance, C-Calib on the AMZ-GOO dataset (Figure 6a) reduced EOD more effectively across all models compared to Calib. Specifically, HierGAT’s EOD for AMZ-GOO was increased to 30.92% from its initial value of 19.50% by Calib; however, C-Calib reduced it to 7.39%. Similarly, on the BEER dataset (Figure 6b), the EO for DITTO and HierGAT is significantly reduced to near-zero values after C-Calib, indicating a strong improvement over Calib. Notably, for DITTO on BEER in Figure 6b, the initial EO was 0.30%, which is already minimal, but C-Calib further reduced it to 0.13%, highlighting the effectiveness of this method. Across other datasets (available in our repository), this pattern of reduced disparity by C-Calib continues, showing that the conditional approach better aligns score distributions across demographic groups by estimating labels.
@@ 6.2.3 Bias Reduction in Conditional Calibration
[…]
@@
6.2.4 Risk of Calibration
To conclude our analysis, we examine how Calib and C-Calib affect performance preservation after calibration. Definition 2 defines risk as the expected absolute change in scores, which directly quantifies how much calibration perturbs the original matcher. In this section, however, we do not directly estimate this score-space risk; instead, we report the change in AUC as an application-facing, threshold-independent proxy for utility. A small AUC change indicates that the calibrated scores largely preserve the ranking quality of the original matcher.
We emphasize that AUC change is not theoretically equivalent to risk in general: small score perturbations can sometimes alter rankings, and conversely some score changes may leave rankings (and hence AUC) unchanged. Thus, our AUC-based analysis should be interpreted as measuring utility impact (performance preservation), complementary to the formal risk notion in Definition 2. Establishing tighter theoretical connections between score deviation and AUC preservation is an interesting direction for future work.
Table 4 shows the effect of Calib and C-Calib on AUC for AMZ-GOO and DBLP-GOO datasets across matching models. For clarity, only the AUC decreases after calibration are reported. In all cases, both algorithms have minimal AUC reductions, and similar results are observed universally, however only a subset of results is reported due to space constraints. In AMZ-GOO dataset, AUC decrease post-calibration using Calib at most 1.01%. Note that a maximum reduction of 1.01% in AUC results in almost completely removing DP bias from an initially significant value. Similarly, in DBLP-GOO dataset, there is almost no change in AUC after calibration using Calib across different matchers, indicating that Calib effectively reduces DP without compromising accuracy.